OpenStudy (anonymous):

What are the solutions of 2x2 - 6x = -8? please help!

OpenStudy (anonymous):

\[2x ^{2}-6x=-8\] This is the equation?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Quadratic equation. First you need to add eight to both sides of the equation which gives you what?

OpenStudy (anonymous):

2x^2-6x+8=0?

OpenStudy (anonymous):

Do you know the quadratic formula?

OpenStudy (anonymous):

Yes but i forget what goes where could you tell me like which is A, b and c?

OpenStudy (anonymous):

Well lets see if we can do it with out the formula. What do we get if we divide both sides of the equation by 2?

OpenStudy (anonymous):

x^2-3x+4=0

OpenStudy (anonymous):

Good now we have to factor this out into (x+a)(x+b)= x^2-3x+4. Can you do this?

OpenStudy (anonymous):

Solve for a and b.

OpenStudy (anonymous):

-2 and -2?

OpenStudy (anonymous):

(x-2)(x-2) does this equal x^2-3x+4? Well x times x equals x^2 and -2 times -2 equals 4. But x times -2 = -2x and -2 times x = -2x leaving us with x^2 -2x-2x+4 or x^2-4x+4. What other factor of 4 can you use?

OpenStudy (anonymous):

-3, -1

OpenStudy (anonymous):

-3 times -1 does not equal 4.

OpenStudy (anonymous):

oh my bad lol -4 -1

OpenStudy (anonymous):

The original problem. All the numbers are correct?

OpenStudy (anonymous):

All the negatives are where there suppose to be?

OpenStudy (anonymous):

I believe so

OpenStudy (anonymous):

Well my error you will have to solve with the quadratic formula. http://ts4.mm.bing.net/th?id=HN.608016830722150363&w=238&h=179&c=7&rs=1&qlt=80&pid=1.7

OpenStudy (anonymous):

Thank you